### v1

# Ginny Wong, Simon See

Abstract
This paper studies the minimal dimension required to embed subset memberships ($m$ elements and $m\choose k$} subsets of at most $k$ elements) into vector spaces, denoted as Minimal Embeddable Dimension (MED). The tight bounds of MED are derived theoretically and supported empirically for various notions of “distances” or “similarities”, including $\ell_{2}$ metric, inner product, and cosine similarity. In addition, we conduct numerical simulation in a more achievable setting, where the $m\choose k$} subset embeddings are chosen as the centroid of the embeddings of the contained elements. Our simulation easily realizes a logarithmic dependency between the MED and the number of elements to embed. These findings imply that embedding-based retrieval limitations stem primarily from learnability challenges, not geometric constraints, guiding future algorithm design.
1 Introduction
Embedding-based retrieval systems answer queries by vector comparison. One common setting of such systems is that the system maintains each element $x_{i}\in X$ as a vector ${\bm{x}}_{i}\in^{d}$, where $X$ is the universe set of elements and d is the vector space [^1] Other metric spaces may apply. d is chosen because it is the most commonly used.. Each query $q$ to answer is also embedded as a vector ${\bm{w}}_{q}\in^{d}$. The answers to a query $q$ are retrieved by (1) comparing ${\bm{w}}_{q}$ against each of the ${\bm{x}}_{i}$ with a scoring function $s(\cdot,\cdot)$ to obtain a score $s_{i|q}=s({\bm{x}}_{i},{\bm{w}}_{q})$ and (2) retrieving the answers with the $k$ largest scores $s_{i|q}$ [^2] Other notions of answer sets may apply, but top-$k$ is chosen also because its common usage.. This paper investigates a fundamental question of the aforementioned retrieval systems. One informal statement is:
Given the universe $X$ with $m$ elements and a scoring function $s$, what is the minimal dimension $d$ of d such that every query with at most $k$ answers is perfectly retrievable?
This question concerns the “Minimal Embeddable Dimension (MED)” $d$, as will be formally defined in Section 2, which depends on the function $s$, the cardinality $m$ of the universe $X$, and also the cardinality $k$ of the answer set to queries we are interested in. For convenience, we say a set of embeddings for elements and subsets is an embeddable configuration if and only if all answers can be retrieved by score comparison.
The study of the minimal embeddable dimension is fundamentally related to several classic research directions, including the VC-dimension in statistical learning theory (Mohri et al., 2018) and the $k$-set problem in combinatorial geometry (Matousek, 2013). It is widely acknowledged that this question is related to the fundamental boundary underlying modern information retrieval (Lee et al., 2019; Weller et al., 2025b), data compression (Andoni and Indyk, 2008), and representation learning (Izacard et al., 2021; Wang et al., 2022).
[…]
Table 1: Tight bounds of minimal dimensions in the standard setting (MED) and the centroid setting (MED-C) of important scoring functions. Theorems are stated in Section 3 and Section 4. $m$ denotes the size of the universe, $k$ indicates the largest size subset to query. Interestingly, the MED does not depend on $m$. The $O(\log m)$ upper bound of MED-C can be easily observed in simulation.MED is the lower bound of MED-C by definition, so the $\Omega(k)$ lower bound of MED-C is omitted in the table.
Scoring function Standard setting (MED) Centroid setting (MED-C)
Lower bound Upper bound Upper bound
Linear (inner product) k1k-1 $2k$ $O(k^{2}\log m)$
Cosine similarity k1k-1 $2k+1$ $O(k^{2}\log m)$
Euclidean distance ($\ell_{2}$) k1k-1 $2k$ $O(k^{2}\log m)$
This paper questions the conclusions by Weller et al. (2025a) by studying the MED problem, as will be formalized in Section 2, from the perspective of approximability. That means we also work on the fundamental property of geometric spaces as in Weller et al. (2025a) but with an extended set of essential scoring functions: in addition to the inner product previously discussed (Weller et al., 2025a), cosine similarity and Euclidean distances are also included due to their wide application in embedding-based systems and representation learning. This paper discusses when those spaces are large enough to contain the subset membership of at most $k$ elements. This perspective, due to its theoretical nature, is not practical enough to help develop scalable ML algorithms. However, it still allows people to determine whether the effectiveness of embedding-based systems is limited by approximability (a fundamental property of geometric spaces) or by learnability (the way we build specific retrieval systems). If it turns out to be an approximability issue, any methodological effort might make little progress before the dimension of the vector space is large enough. Otherwise, there is still hope in developing more performant embedding-based retrieval systems.
Theoretical Contribution. We establish a clear quantitative relation $d=\Theta(k)$ between the MED $d$ and the maximum cardinality $k$ of the answer set for all three scoring functions. The major results are presented in Table 1, “Standard setting (MED)”. Our theoretical results are independent of the universe’s size $m$, suggesting that the fundamental property of geometric space does not limit the embedding-based retrieval system. They sharpen the theoretical bounds by Weller et al. (2025a) and contradict the empirical results in Weller et al. (2025a).
Empirical Support. We also conduct numerical simulations to show that the minimal dimension, where an embeddable configuration of embeddings can be found by optimization, does not grow polynomially with $m$, but at most logarithmically. The numerical simulations also justify the correctness of our tight bounds in MED and credit the limitation of the effectiveness of embedding-based systems to learnability rather than approximability. Some brief details are presented here, and more information can be found in Section 4.
In our simulation, we further assume the query vector of $q$ to be given by ${\bm{w}}_{q}:=\frac{1}{|S|}\sum_{x\in S}{\bm{x}}$ as the centroid of vectors in the answer set $S$ and denoted as the centroid setting. Our centroid setting is “more achievable” than the “free embedding optimization” setting (Weller et al., 2025a) because our setting only requires optimising $m$ embeddings for elements, but requires no optimization of the $\binom{m}{k}$ embeddings of the subsets. Simulations in both settings only show the existence of an embeddable configuration of $m$ elements in specific dimensions. By definition, they reveal upper bounds on MED.
Another interesting finding is that the simulation in the centroid setting reveals an empirical logarithmic upper bound of MED. This $O(\log m)$ dependency agrees with our theoretical upper bounds from a probabilistic construction. The minimal embeddable dimension in the centroid setting is denoted MED-C, and listed in Table 1. Compared to the polynomial dependency revealed by the free embedding optimization (Weller et al., 2025a), it might be counterintuitive because the centroid setting has fewer degrees of freedom ($\binom{m}{k}$ subset query embeddings are not optimized in the centroid setting), providing a perfect example of the bottleneck in learnability rather than approximability. More discussion can be found in Section 4.
In summary, our work reframes the debate from geometric limits to learnability, offering both theoretical guarantees and empirical evidence that low-dimensional embeddings suffice for perfect retrieval. This paper is organized as follows. Section 2 formally defines the Minimal Embeddable Dimension (MED), establishes some inequalities of MED, and also introduces the centroid setting. Section 3 proves the tight bounds of the MED under various scoring function classes. Section 4 derives the bounds for MED in centroid settings and validates them through numerical simulations. Section 5 discusses two highly related works. Finally, Section 6 concludes the paper with limitations and future directions.
2 Minimal Embeddable Dimension
For convenience, we summarize the notations in this paper. $m,n,d,k$ are positive integers. $n$ and $d$ are both used for dimension. They can be used interchangeably. The subtle difference is that $d$ usually represents dimension in general while $n$ appeared in the quantitative relations. $X$ is the set of $m$ elements to be embedded and ${\bm{x}}\in^{d}$ denotes the embeddings of $x\in X$. For simplicity, we don’t distinguish the embedding $\bm{x}$} and element $x$, so it is fine to state $X=\{{\bm{x}}_{i}\}_{i=1}^{m}\subseteq^{d}$. We study all possible queries $q$ that concern at most $k$ objects in $X$. In that sense, it is equivalent to consider all subsets of at most $k$ elements in $X$ ($k\leq m$ by default), denoted as $\mathcal{C}_{k}=\{S\subseteq X,|S|\leq k\}$ and $\mathcal{C}_{m}=2^{X}$. We also use $q$ to denote a subset in $X$. The scoring function $s:^{d}\times^{d}\mapsto$ measures the relatedness of two vectors in d. The description of d is omitted if the context is clear. The scoring functions of our interest include:
Linear:
$s_{\rm linear}({\bm{x}},{\bm{w}})=\langle{\bm{x}},{\bm{w}}\rangle$, where $\bm{w}$} is a query vector, $\langle\cdot,\cdot\rangle$ is the inner product.
Cosine similarity:
$s_{\cos}({\bm{x}},{\bm{w}})=\frac{\langle{\bm{x}},{\bm{w}}\rangle}{\|{\bm{x}}\|\|{\bm{w}}\|}$.
Euclidean distance ($\ell_{2}$):
$s_{\ell_{2}}({\bm{x}},{\bm{w}})=-\|{\bm{x}}-{\bm{w}}\|_{2}$.
[…]
The primary focus of this section is to present the definition and general properties of the Minimal Embeddable Dimension (MED). Notably, we dedicate the last subsection 2.3 to a special centroid setting, where the ${\bm{w}}_{q}=\frac{1}{|S|}\sum_{x\in S}{\bm{x}}$, which concerns the MED in centroid (MED-C), an upper bound of MED.
2.1 $k$-shatter problem
To formally define the minimal embeddable dimension, we introduce the concept of $k$-shattering.
Definition 2.1 ($k$-shattering).
Let $X\subseteq^{d}$ be a set of $m$ points. $X$ is $k$-shattered by $\mathcal{F}$ if and only if $\forall S\in\mathcal{C}_{k},\exists f_{S}\in\mathcal{F},\forall{\bm{x}}\in S,\forall{\bm{y}}\notin S,f_{S}({\bm{x}})>b_{S}>f_{S}({\bm{y}})$, where $b_{S}\in$ depends on $S$ and $\mathcal{F}$.
Remark 2.2.
The definition of $k$-shattering precisely determines whether there exists a configuration of vectors in $X$ under a specific functional family or scoring function with query embeddings such that the embedding-based retrieval built upon this configuration succeeds on all queries concerning at most $k$ elements.
Minimal Embeddable Dimension (MED) is then defined based on $k$-shattering, which depends on $m$, $k$, and $\mathcal{F}$. For convenience, MED is denoted as a function $n^{*}=\textsc{MED}(m,k;\mathcal{F})$.
Definition 2.3 (Minimal Embeddable Dimension).
Given $m$, $k$, $\mathcal{F}$, MED $n^{*}$ is the integer that a configuration of $m$ points that can be $k$-shattered by $\mathcal{F}$ exists in $^{n^{*}}$ but not in $^{(n^{*}-1)}$.
One direct result, according to Definition 2.3, is the non-strict monotonicity of $\textsc{MED}(m,k;\mathcal{F})$.
Proposition 2.4.
For $2\leq k\leq m$, the following inequality holds:
$\textsc{MED}(m,k-1;\mathcal{F})$ $\leq\textsc{MED}(m,k;\mathcal{F})$
$\leq\textsc{MED}(m+1,k;\mathcal{F}).$
Meanwhile, when all subsets of at most half the points can be shattered, all subsets can be shattered.
Proposition 2.5.
For $k\geq\lfloor\frac{m}{2}\rfloor$, $\textsc{MED}(m,k;\mathcal{F})=\textsc{MED}(m,m;\mathcal{F})$.
2.2 General bounds of MED by VC dimension
The definition of $k$-shattering also defines the VC dimension (Mohri et al., 2018), which is redefined below in terms of $k$-shattering.
Definition 2.6 (VC dimension).
The VC dimension of a functional family $\mathcal{F}$ (maps n to ) is the maximal size $m$ of a set $X$ that can be $m$-shattered by $\mathcal{F}$, denoted as $m=\textsc{VCD}(n;\mathcal{F})$.
VC dimension is a measure of the capacity of classes of sets or binary functions (Vapnik, 2013). This concept plays a fundamental role in statistical learning theory, particularly in the study of approximability. By defining both the MED and VC dimensions using $k$-shattering, it is reasonable to expect some connections to hold between them, and indeed they do. For notational convenience, we define the “inversion" of VC dimension as
$\textsc{VCD}^{-1}(m;\mathcal{F})=\min\{i\mid s.t.\textsc{VCD}(i;\mathcal{F})\geq m\}$
,namely the minimum dimension required. Notably, the function $\textsc{VCD}(;\mathcal{F})$ is not guaranteed to increase strictly; therefore, there is no inverse function formally.
Lemma 2.7.
$\textsc{MED}(m,m;\mathcal{F})=\textsc{VCD}^{-1}(m;\mathcal{F})$.
Proof.
Let $n=\textsc{VCD}^{-1}(m,\mathcal{F})$, then (1) $m$ points in n can be $m$-shattered by $\mathcal{F}$ but (2) $m$ points in $^{(}n-1)$ cannot be $m$-shattered by $\mathcal{F}$. Those two claims imply, in the sense of MED, that (1) $\textsc{MED}(m,m,\mathcal{F})\leq n$ and (2) $\textsc{MED}(m,m,\mathcal{F})>n-1$. Thus $\textsc{MED}(m,m,\mathcal{F})=n$
Additionally, combining Proposition 2.4 and Lemma 2.7 yields the following proposition.
Proposition 2.8.
$\textsc{VCD}^{-1}(k;\mathcal{F})\leq\textsc{MED}(m,k;\mathcal{F})\leq\textsc{VCD}^{-1}(m;\mathcal{F})$.
Thus, a rough lower and upper bound of MED can be derived from the VC dimension. Specifically, the upper (lower) bounds of VC dimensions now form the lower (upper) bounds of MED, respectively.
2.3 MED in the centroid setting
We present the centroid setting, where the query embedding is the simple average of the element embeddings. The centroid setting imposes more constraints than $k$-shattering. Formal definitions are now provided.
We begin with the definition of $k$-centroid shattering as the restricted version of $k$-shattering.
Definition 2.9 ($k$-centroid shattering).
Let $X\subseteq^{d}$ be a set of $m$ points, $X$ is $k$-centroid shattered by a scoring function $s(\cdot,\cdot)$ if and only if $\forall S\in\mathcal{C}_{k},\forall{\bm{x}}\in S,\forall{\bm{y}}\notin S,s({\bm{x}},{\bm{c}}_{S})>s({\bm{y}},{\bm{c}}_{S})$, where ${\bm{c}}_{S}=\sum_{x\in S}\frac{1}{|S|}{\bm{x}}$ is the center vector of $S$.
Remark 2.10.
We stress that the concept of centroid shattering concerns a specific scoring function ($s_{\rm linear},s_{\cos},s_{\ell_{p}}$) rather than a functional class $\mathcal{F}_{\rm linear},\mathcal{F}_{\cos},\mathcal{F}_{\ell_{p}}$ because we already fix the vector for each of the query as the center of its answer set $S$, specifically, ${\bm{w}}:={\bm{c}}_{S}$.
Then, the MED in the centroid setting is defined below,
Definition 2.11 (MED in Centroids (MED-C)).
Given $m,k,s(\cdot,\cdot)$, MED-C $n^{*}=\textsc{MED-C}(m,k;s)$ is the integer that a configuration of $m$ points that can be $k$-centroid shattered by $s(\cdot,\cdot)$ exists in $^{n^{*}}$ but not in $^{(n^{*}-1)}$.
MED-C and MED. In the original definition of MED, only the existence of functionals is required to prove the $k$-shattering definition. In the centroid setting, $k$-centroid shattering fixes the subset query embeddings ${\bm{c}}_{S}$, which further determine functionals explicitly. Noticing that the number of “free” functionals involved in $k$-shattering is $\binom{m}{k}$, the $k$-centroid setting reduces the freedom into as few as $m$ vectors, which facilitates numerical simulation. For the same reason, however, MED lower bounds MED-C because of the flexibility to freely determine $f(\cdot)$ other than the centroid one $s(\cdot,{\bm{c}}_{S})$.. Formally, we have the following proposition.
Proposition 2.12.
$\textsc{MED}(m,k;\mathcal{F})\leq\textsc{MED-C}(m,k;s)$, where $\mathcal{F}$ is the functional family induced by the scoring function $s$.
Proof.
Let $n=\textsc{MED-C}(m,k;s)$, we consider the configuration of vectors ${\bm{x}}_{i}$ for i=1,,mi=1,...,m elements. Let the functionals for each $S\in\mathcal{C}_{k}$, we select the functional $f_{S}(\cdot):=s(\cdot,{\bm{c}}_{S})$, where ${\bm{c}}_{S}=\frac{1}{|S|}\sum_{x\in S}{\bm{x}}$. We can say that this configuration is also $k$-shattered by $\mathcal{F}$. Then, $\textsc{MED}(m,k;\mathcal{F})\leq\textsc{MED-C}(m,k;s)$ by definition. ∎
MED-C and neural set embeddings. One of the common deep learning approaches to compute a set embedding is to aggregate the embeddings of contained elements (Zaheer et al., 2017). We can also study the Minimum Embeddable Dimension in the Neural set embedding setting (MED-N), where the set embedding ${\bm{w}}_{S}={\rm NN}(\{{\bm{x}}:x\in S\})$ is computed by a complex neural architecture (such as MLP or multi-head self attention). The centroid setting simplifies the aggregation of complex neural networks by simple averaging. Therefore, one could expect that the MED-C is not smaller than the MED-N due to a less capable averaging operation. Combined with the definition of MED, this informal discussion suggested that we can roughly consider MED-N to be between MED and MED-C.
3 Tight bounds for MED and discussions
This section derives tight bounds of MED under $\mathcal{F}_{\rm linear}$, $\mathcal{F}_{\cos}$, and $\mathcal{F}_{\ell_{2}}$. The choices are based on the broad interest of their applications in modern machine learning and representation learning. By Proposition 2.5, we only need to study the cases where $2\leq k\leq\lfloor\frac{m}{2}\rfloor$. The general proof strategy applied here is to find the upper bound of MED by construction, and to find the lower bound by VC dimension by Proposition 2.8.
3.1 Tight bounds of $\textsc{MED}(m,k;\mathcal{F}_{\rm linear})$
To begin with, let’s consider the cyclic polytope (Ziegler, 2012).
Example 3.1 (Cyclic polytope).
Given moment curve ${\bm{x}}(t)=(1,t,t^{2},...,t^{d})\in^{d},0\leq t\leq 1$, a cyclic polytope is the convex hull ${\rm Conv}({\bm{x}}(t_{1}),{\bm{x}}(t_{2}),\dots,{\bm{x}}(t_{m}))$, where $t_{i},i=1,...,m$ can be arbitrary real number.
One of the most well-known results of cyclic polytopes is that a cyclic polytope in d is an $\lfloor\frac{d}{2}\rfloor$-neighborly polytope, in which every $k$ vertices form a face when $k\leq\lfloor\frac{d}{2}\rfloor$ (Ziegler, 2012). By forming a $k$-face, it means that the $k$ vertices in this face can be linearly separated from the rest of the vertices of the polytope.
Theorem 3.2.
$k-1\leq\textsc{MED}(m,k;\mathcal{F}_{\rm linear})\leq 2k.$
Proof.
Noticing the VC dimension of $\mathcal{F}_{\rm linear}$ in n is $n+1$ (Mohri et al., 2018), the lower bound is derived as a direct result of Proposition 2.8. The upper bound is by the cyclic polytope construct in Example 3.1. ∎
We see that $\textsc{MED}(m,k,\mathcal{F}_{\rm linear})$ only depends on $k$. Ignoring the coefficient, $\textsc{MED}(m,k;\mathcal{F}_{\rm linear})=\Theta(k)$ also holds. Then, we show that the $\mathcal{F}_{\cos}$ and $\mathcal{F}_{\ell_{2}}$ share the similar bounds as $\mathcal{F}_{\rm linear}$.
3.2 Tight bounds of $\textsc{MED}(m,k;\mathcal{F}_{\ell_{2}})$
By geometric constructions in n regarding the $k$-shattering, the following relation is revealed.
Proposition 3.3.
$\textsc{MED}(m,k;\mathcal{F}_{\ell_{2}})\leq\textsc{MED}(m,k;\mathcal{F}_{\rm linear}).$
Proof.
Given a configuration that can be $k$-shattered by $\mathcal{F}_{\rm linear}$, then for each set $S,|S|\leq k$, there exists a hyperplane $H_{S}$ that separates $S$ from XSX-S. We can always find an $\ell_{2}$ ball that includes $S$ and is tangent to $H_{S}$, which automatically does not contain XSX-S. ∎
When considering the Proposition 3.3, we conclude $\textsc{MED}(m,k;\mathcal{F}_{\ell_{2}})$ in Theorem 3.4 with additional information that $\textsc{VCD}(n;\mathcal{F}_{\ell_{2}})=n+1$ and Proposition 2.8.
Theorem 3.4.
$k-1\leq\textsc{MED}(m,k;\mathcal{F}_{\ell_{2}})\leq 2k.$
3.3 Tight bounds of $\textsc{MED}(m,k;\mathcal{F}_{\cos})$
Proposition 3.5.
$\textsc{MED}(m,k;\mathcal{F}_{\rm linear}))$ $\leq\textsc{MED}(m,k;\mathcal{F}_{\cos})$
$\leq\textsc{MED}(m,k;\mathcal{F}_{\rm linear}))+1.$
Proof sketch: Notice that the decision boundary by cosine similarity on a sphere is the intersection between hyperplanes and this sphere. Given a configuration in n that can be $k$-shattered by $\mathcal{F}_{\rm linear}$, one could use the inverse stereographic projection to project onto a sphere in n+1, where the original points are projected to the points on the sphere. Meanwhile, suppose a configuration can be $k$-shattered by cosine similarities. In that case, one can first project every point onto the sphere by $x\mapsto\frac{x}{\|x\|_{2}}$; then, the decision boundaries on spheres can be extended to hyperplanes. The full proof can be found in Appendix A.1.
Combination of the Theorem 3.2 and the Proposition 3.5 also shows that $\textsc{MED}(m,k;\mathcal{F}_{\cos})=\Theta(k)$.
Theorem 3.6.
$k-1\leq\textsc{MED}(m,k;\mathcal{F}_{\cos})\leq 2k+1$.
3.4 Discussion on $\Theta(k)$ bound of MEDs
To summarize, the minimum embeddable dimension for $\mathcal{F}_{\rm linear}$, $\mathcal{F}_{\cos}$, and $\mathcal{F}_{\ell_{2}}$ are all $\Theta(k)$, which is independent of the number of total elements in the set system. This result suggests that, despite the difficulty, it is possible to encode all top-$k$ queries into a space of no more than $2k$ dimensions, regardless of the number of elements to embed.
The real performance bottleneck of the embedding-based retrieval. The bounds clearly show that the geometric space does not limit the effectiveness of the embedding-based retrieval system. To make this even more transparent, one possible strategy to build a perfect top-$k$ queries of a retrieval system with $m$ elements is to (1) put elements in the vertices of a cyclic polytope in 2k, (2) learn the query embeddings model which predict the query embedding that clearly separates the (at most $k$) answers from the rest. Our theory guarantees that such query embeddings exist, but the real problem is how to build a function to predict them. Knowing that neural networks can universally approximate functions (Hornik et al., 1989), the bottleneck of embedding-based retrieval is essentially a learnability problem: how to actually learn such a function with embedding dimension $2k$ from data.
Can the $\Theta(k)$ bounds be observed? Although the fundamental property of the underlying geometric space does not put any restrictions on embedding-based retrieval. However, two issues still challenge whether we can observe this bound in the real world. (1) limited numerical precision: Even with the cyclic polytope example, it is still questionable when $m$ embeddings are represented in a fixed-digit floating-point number. The budget of float points does not have infinite capacity. (2) infeasible number of functionals to determine: To realize $\Theta(k)$ bound, one would search for $\binom{m}{k}$ functionals. The total number of tasks is infeasible to complete. Those two issues
4 Upper bounds of MED-C, Simulation, and Implications
This section discusses the minimal embeddable dimension in the centroid setting (MED-C). We will see that MED-C grows logarithmically with the universe’s size $m$ in both theory and experiments.
4.1 Upper bounds
For inner product, cosine similarity, and Euclidean distance, we use the probabilistic method as a common proving strategy to prove the logarithmic upper bound. In short, we sample $m$ element embeddings from a probabilistic distribution in n and find the conditions on $k$, $m$, and $n$ such that the desired $k$-centroid shattering configuration occurs with positive probability. The random vectors here are part of the probabilistic construction and do not directly link to the simulation.
Theorem 4.1.
$\textsc{MED-C}(m,k;s_{\rm linear})=O(k^{2}\log m)$
Proof sketch: We consider the probabilistic method, where $v_{1},\dots,v_{m}\sim\mathcal{N}(0,1/n)$ in n, we show there is positive probability such that for any subset $S$, $|S|\leq k$, $v_{1}\in S$ and $v_{2}\notin S$ $\langle v_{1},\sum_{u\in S}u\rangle>\langle v_{2},\sum_{u\in S}u\rangle$ under the condition of $n>Ck^{2}\log m$. The full proof can be found in Appendix A.2.
Following a similar probabilistic construction, we can also prove that $\mathcal{F}_{\cos}$ and $\mathcal{F}_{\ell_{2}}$ share the same upper bound. The proofs are also in the Appendix A.2.
Theorem 4.2.
$\textsc{MED-C}(m,k;\mathcal{F}_{\cos})=O(k^{2}\log m).$
Theorem 4.3.
$\textsc{MED-C}(m,k;\mathcal{F}_{\ell_{2}})=O(k^{2}\log m).$
Although the $O(k^{2}\log m)$ bound may not be tight, it still suffices to make significant implications. An easy lower bound for MED-C is the lower bound of MED, which is $\Omega(k)$. We can see that the gap between the lower and upper bounds is a factor of $O(\log m)\times O(k)$. We discuss those two factors separately. For the factor of $O(\log m)$, it still demonstrates that a sublinear growth for MED-C with the size $m$ of the universe is feasible. For more complex neural set embedding settings, such as MED-N discussed in Section 2.3, a similar upper bound also applies. For the additional factor $O(k)$, we would like to convince the readers that $O(k\log m)$ is not very easy to achieve, so the $O(k^{2}\log m)$ is not very bad. To the best of the authors’ knowledge, the most similar (but not exact) problem setting to MED-C is 1-bit compress sensing (Aksoylar and Saligrama, 2014), with a lower bound of $\Omega(k\log\frac{m}{k})$, which cancels the additional $O(k)$ factor (compare $O(k^{2}\log m)$ for MED-C with $\Theta(k)$ for MED) but requires an additional sparse recovery algorithm rather than simply comparing scores independently. This related work suggests that the extra cost should be paid to achieve the bound $O(k\log m)$ of MED-C.
4.2 Numerical simulation
The situation where $k\ll m$ is of particular interest because there are many documents to be retrieved, and only a very small subset is required in a real-world scenario. Therefore, numerical simulation typically focuses on large $m$ and small $k$ to demonstrate the effectiveness of the bounds. In the numerical simulation, we use "critical" to denote the largest possible number of objects for a given dimension, or the smallest dimension required to accommodate a specified number of objects.
Experiment settings. We use optimize $m$ embeddings randomly initialized in by standard normal distribution in d. The objective function is a hinge loss computed over all pairwise positive and negative pairs across the top-$k$ queries. We use the Adam optimizer (Kingma, 2014) with a learning rate of 1. The optimization stops either after 1,000 steps or after every query finds its perfect answers. To facilitate full reproducibility, we list our Python code in the Appendix to disclose other details.
Comparison with baselines Previous work (Weller et al., 2025a) didn’t provide a quantitative relation in their theory, but instead offered an empirical relation via optimization of free embeddings, which corresponds to the standard setting in this paper. They search for the critical cardinality of the universe $m^{*}$ in a fixed dimension $d$ for the $k=2$ case, and established the following empirical results. We list their curve in Equation 1.
$m_{\rm WBNL}(d)=-10.5322+4.0309d+0.0520d^{2}+0.0037d^{3}.$ (1)
This result yields a pessimistic implication: the largest number of objects we can support grows with $d$ only cubically. However, our bounds in MED-C suggests that the number of objects grows exponentially with $d$.
It is a crucial question: how can the critical points identified by simulation in the centroid setting be compared with those found in the free embedding optimization setting (Weller et al., 2025a). The goal of numerical simulation is to estimate the minimum dimension such that $m$ elements can be $k$-shattered in some sense (or the maximum number of elements can be located in a given dimension). Therefore, all critical points found by any numerical simulations only suggest an upper bound of the minimal dimension (or lower bound of the maximum number of elements). The goal of this paper is to show that the MED (or the lower bound) grows slowly enough so that the fundamental property of the geometric space does not bottleneck the embedding-based retrieval. Therefore, results from simulations, whether computed from a centroid setting or a free embedding optimization setting, can serve as empirical evidence of the MED upper bound.
To compare with their empirical results, we also ran the experiments with $k=2$ on the centroid setting. We plot the results in Figure 1(a) and Figure 1(b). Figure 1(a) suggests that the number of critical points found in our centroid setting surpasses the curve fitted by Weller et al. (2025a) easily. From Fig 1(b), we can see from the results that our theory of MED-C agrees well with the empirical log-linear fitting, which is of a lower complexity family than the fitted result in Weller et al. (2025a).
Figure 1: (a) The comparison of the growth of the critical number of points in our simulation and the curve fitted in Equation 1.; (b) The comparison of the growth of the critical dimensions in our simulation and the curve fitted in Equation 1, and the $x$ axis is plotted in a log scale.
4.3 Practical implication
Regarding the bottleneck of the embedding-based retrieval. The empirical results above suggested that the actual limitation of embedding-based retrieval, i.e., the number of objects that can be embedded, actually grows at least exponentially with the dimension. This contradicts the previous claims made by Weller et al. (2025a): The stagnation of the performance in the embedding-based system does NOT necessarily come from the fundamental geometric property of the vector space. The empirical findings resonated with the conclusion from $\Theta(k)$ bounds in a weaker form: we can observe the exponential growth of critical points and logarithmic growth of MED, but not a vertical line or horizontal line.
Why numerical simulation in centroid setting achieves stronger points than free embedding optimization? It is also a little bit counterintuitive that, optimizing with $\Theta(\binom{m}{k})$, more degrees of freedom actually leads to weaker results. Specifically, one should expect the optimums of the free embedding optimization and centroid setting to follow the following inequality:
$\overbrace{\min_{{\bm{w}}_{S}:S\in\mathcal{C}_{k}}}^{\text{additional variables}}\min_{{\bm{x}}:x\in X}L_{\rm FreeEmbedding}\leq\min_{{\bm{x}}:x\in X}L_{\rm Centroid},$ (2)
where we define hinge losses $L_{\rm FreeEmbedding}:=\sum_{S\in\mathcal{C}_{k},x\in S,y\notin S}\max(0,s({\bm{y}},{\bm{w}}_{S})-s({\bm{x}},{\bm{w}}_{S}))$ and $L_{\rm Centroid}:=\sum_{S\in\mathcal{C}_{k},x\in S,y\notin S}\max(0,s({\bm{y}},\frac{1}{|S|}\sum_{x\in S}{\bm{x}})-s({\bm{x}},\frac{1}{|S|}\sum_{x\in S}{\bm{x}}))$ are optimization objectives in free-embedding setting and centroid setting, respectively.
However, the simulation empirically reveals the inequality in the other direction. It is another perfect example of how optimization affects the final results and further distorts the empirical findings.
5 Related Theoretical Works
Given the fundamental nature of embedding-based retrieval, we omit a survey of empirical efforts and instead discuss related theoretical contributions. In particular, two works considered related settings are also worth discussing.
Guo et al. (2019) established the dependency of dimension $d$ and $K$ (the total number of queries) and $s$ (maximal number of queries associated with each object), as $d=O(s^{2}\log(K))$ in their Theorem 2.1. Applying to our study of MED, the total number of top $k$ queries of $m$ objects is $K:=\binom{m}{k}$, and the maximal number of queries on one object is $s:=\binom{m-1}{k-1}$. Plugging those two values back suggests an upper bound of MED by $O(\binom{m-1}{k-1}^{2}\log\binom{m}{k})$, which is worse than our $\Theta(k)$ bound.
You et al. (2025) studies the embedding-based retrieval in a specified "Hierarchical Retrieval (HR)" setting. It assumes a directed acyclic graph $\mathcal{G}$ between objects. The answer set $S(q_{i})$ of one query $q_{i}$ is defined by all "reachable objects" in $\mathcal{G}$ from the exact match object $E(q_{i})$. The upper bound is established as $d=O(\max(s\log m,\frac{1}{\epsilon^{2}}\log m)$, where $s$ is the largest cardinality of all answer sets. Applying their result in our top-$k$ query case, the graph $G$ is a $k$-clique. It still suggested a $\max(O(k\log m),O(\frac{1}{\epsilon^{2}})\log m)$ upper bound of MED, which is worse than our $\Theta(k)$ bound.
6 Conclusion and future work
This paper establishes the theoretical bounds for the minimal embeddable dimensions and studies an additional centroid setting. It revealed a positive fact: we can use a small number of dimensions to support an extensive retrieval system, provided we consider only queries with small cardinality. Meanwhile, we can clearly conclude that the effectiveness of embedding-based retrieval systems is limited by learnability (the way we build the specific retrieval system) and debunk the claim from the approximability that the theoretical limitation lies in the fundamental property of the vector space that prevents the subset membership structure from being approximated.
Future work should include more discussion of advanced embedding spaces, such as hyperbolic and Wasserstein spaces; a more sophisticated structure for queries and their answers, including situations where the cardinality of answer sets follows a power-law distribution; and a more realistic setting of fixed-point float numbers, such as float point 32, or even float point 4.
Supplementary Material
Appendix A Additional Proofs
A.1 Proof of Proposition 3.5
Proof.
Let $n=\textsc{MED}(m,k;\mathcal{F}_{\cos})$ and there exists $X=\{x_{1},\dots,x_{m}\}\subset\mathbb{R}^{n}\setminus\{0\}$  [^4] If there is $x_{i}=0$, the proof is basically the same, we omit this situation for convenience.that is $k$-shattered by $\mathcal{F}_{\cos}$. Define the radial projection $\rho:\mathbb{R}^{n}\setminus\{0\}\to S^{n-1}$ by $\rho(x)=\frac{x}{\|x\|_{2}}$, and let $Y=\{\rho(x_{i})\}_{i=1}^{m}\subset S^{n-1}$.
By $k$-shattering, for each $S\subseteq X$ with $|S|\leq k$ there exist ${\bm{w}}_{S}\in\mathbb{R}^{n}\setminus\{0\}$ such that
$x\in S\Rightarrow r_{S}-\frac{\langle{\bm{w}}_{S},x\rangle}{\|{\bm{w}}_{S}\|\|x\|}\leq 0,\qquad x\in X\setminus S\Rightarrow r_{S}-\frac{\langle{\bm{w}}_{S},x\rangle}{\|{\bm{w}}_{S}\|\|x\|}>0.$
For $z\in S^{n-1}$ define the affine functional
$f_{S}(z)\;:=\;\big\langle\tfrac{-{\bm{w}}_{S}}{\|{\bm{w}}_{S}\|},\,z\big\rangle+r_{S}\ \in\ \mathcal{F}_{\rm linear}.$
Then, for every $x\in X$ we have
$f_{S}(\rho(x))\;=\;r_{S}-\frac{\langle{\bm{w}}_{S},x\rangle}{\|{\bm{w}}_{S}\|\|x\|},$
so the inequalities above becomes
$y\in\rho(S)\Rightarrow f_{S}(y)\leq 0,\qquad y\in Y\setminus\rho(S)\Rightarrow f_{S}(y)>0.$
Therefore we show that
$\textsc{MED}(m,k;\mathcal{F}_{\rm linear})<=\textsc{MED}(m,k;\mathcal{F}_{\cos})$
Conversely, let $n^{\star}=\mathrm{MED}(m,k;\mathcal{F}_{\rm linear})$. Then there exists $X=\{x_{1},\dots,x_{m}\}\subset\mathbb{R}^{n^{\star}}$ such that for every $S\subseteq X$ with $|S|\leq k$ there is $f_{S}(x)=\langle{\bm{w}}_{S},x\rangle+b_{S}$ satisfying
$x\in S\Rightarrow f_{S}(x)\leq 0,\qquad x\in X\setminus S\Rightarrow f_{S}(x)>0.$
Consider the embedding $\phi:\mathbb{R}^{n^{\star}}\to S^{n^{\star}}\subset\mathbb{R}^{n^{\star}+1}$ given by
$\phi(x)\;=\;\frac{(x,1)}{\|(x,1)\|},$
where $(x,1)$ denotes the concatenation of $x$ and 1.
Let $Y=\{\phi(x_{i})\}_{i=1}^{m}\subset S^{n^{\star}}$. For each $S$, set
$u_{S}\;=\;\frac{({\bm{w}}_{S},b_{S})}{\|({\bm{w}}_{S},b_{S})\|}\in S^{n^{\star}},\qquad t_{S}:=0,$
and define $g_{S}(y)=\tfrac{\langle u_{S},y\rangle}{\|u_{S}\|\|y\|}-t_{S}=\langle u_{S},y\rangle$ (since $\|u_{S}\|=\|y\|=1$ on $S^{n^{\star}}$). Then for every $x\in\mathbb{R}^{n^{\star}}$,
$g_{S}(\phi(x))=\left\langle\frac{({\bm{w}}_{S},b_{S})}{\|({\bm{w}}_{S},b_{S})\|},\,\frac{(x,1)}{\|(x,1)\|}\right\rangle=\frac{\langle{\bm{w}}_{S},x\rangle+b_{S}}{\|({\bm{w}}_{S},b_{S})\|\cdot\|(x,1)\|}.$
Therefore, for $x\in X$ we have the exact shattering
$f_{S}(x)\leq 0\iff g_{S}(\phi(x))\leq 0,\qquad f_{S}(x)>0\iff g_{S}(\phi(x))>0.$
A.2 Proof of Theorem 4.1, Theorem 4.2, and Theorem 4.3
Those three proofs are very similar, so they are grouped together.
Proof.
Let $v_{1},\dots,v_{m}\sim\mathcal{N}(0,I_{n}/n)$ in n, our goals are to show that there is a positive probability such that for any subset $S$, $|S|\leq k$, $v_{1}\in S$ and $v_{2}\notin S$, the following inequalities holds:
[…]
To show this, considering the concentration of the inner product of two independent random vectors $v_{i}$ and $v_{j}$ drawn from $\mathcal{N}(0,\frac{1}{n})$:
$\Pr\left(|\langle v_{i},v_{j}\rangle|\geq\frac{1}{3k}\right)\leq 2\exp\left(-c\frac{n}{k^{2}}\right),$ (6)
where $c$ is a constant.
Also, noticing that the norm of such vectors is concentrated as
$\Pr\left(|\|v_{i}\|-1|\geq\frac{1}{3k}\right)\leq 2\exp\left(-c\frac{n}{k^{2}}\right),$ (7)
To prove theorem 4.1, let $k=l$. For $m$ vectors, take the union bound of (1) inner products of all possible pairs and (2) the norm of all vectors, then we derive the probability that any of the inner products (or the $\|v_{i}-1\|$) is greater than $\frac{1}{3k}$, and then force it to be smaller than 1.
$2\left(m+\frac{m(m-1)}{2}\right)\exp\left(-c\frac{n}{k^{2}}\right)<1.$ (8)
The probability of the event where any inner products are smaller than $\frac{1}{3k}$, and norms are no diverge from 1 larger than $\frac{1}{3k}$ is positive when
$n>Ck^{2}\log m.$ (9)
where $C$ is a constant.
Under such conditions,
$LHS=\langle v_{1},\sum_{u\in S}u\rangle=\|v_{1}\|+\sum_{u\in S-\{v_{1}\}}\langle v_{1},u\rangle>1-\frac{1}{3k}+(|S|-1)\frac{1}{3k}=1-\frac{|S|}{3k},$ (10)
$RHS=\sum_{u\in S}\langle v_{2},u\rangle>\frac{|S|}{3k}.$ (11)
Because $|S|\leq k$, so $LHS>RHS$, Theorem 4.1 is proved.
To prove Theorem 4.2
$LHS$ $=\langle\frac{v_{1}}{\|v_{1}\|},\sum_{u\in S}u\rangle>\frac{1-|S|/3k}{1+1/3k}=\frac{3k-|S|}{3k+1}$ (12)
$RHS$ $=\langle\frac{v_{2}}{\|v_{2}\|},\sum_{u\in S}u\rangle<\frac{|S|}{3k(1-\frac{1}{3k})}=\frac{|S|}{3k-1}.$ (13)
Then,
$LHS-RHS>\frac{3k(3k-2|S|-1)}{9k^{2}-1}.$ (14)
Because $1<|S|\leq k$, so LHSRHS>0LHS-RHS>0 and Theorem 4.2 is proved.
To prove Theorem 4.3
$LHS$ $=\|v_{1}-\frac{1}{|S|}\sum_{u\in S}u\|_{2}^{2}=\|v_{1}\|_{2}^{2}+\|\frac{1}{|S|}\sum_{u\in S}u\|_{2}^{2}-\frac{2}{|S|}\langle v_{1},\sum_{u\in S}u\rangle$ (15)
$<1+\frac{1}{3k}-\frac{2}{|S|}(1-\frac{|S|}{3k})+\|\frac{1}{|S|}\sum_{u\in S}u\|_{2}^{2},$ (16)
$RHS$ $=\|v_{2}-\frac{1}{|S|}\sum_{u\in S}u\|_{2}^{2}=\|v_{2}\|_{2}^{2}+\|\frac{1}{|S|}\sum_{u\in S}u\|_{2}^{2}-\frac{2}{|S|}\langle v_{2},\sum_{u\in S}u\rangle$ (17)
$<1-\frac{1}{3k}-\frac{2}{|S|}\frac{|S|}{3k}+\|\frac{1}{|S|}\sum_{u\in S}u\|_{2}^{2}.$ (18)
Then,
$LHS-RHS<\frac{2}{3k}+\frac{2}{3k}-\frac{2}{|S|}(1-\frac{|S|}{3k})=\frac{2}{k}-\frac{2}{|S|}.$ (19)
Because $|S|\leq k$, so LHSRHS<0LHS-RHS<0 and Theorem 4.3 is proved. ∎
Appendix B Simulation Code
1import torch
2from torch import optim
3from tqdm import trange
4import itertools
5import matplotlib.pyplot as plt
6import numpy as np
7
8class Trainer:
9 def __init__(self, n, k):
10 self.n = n
11 self.k = k
12 self.device="cuda" if torch.cuda.is_available() else "cpu"
13 print("Usingdevice", self.device)
14 self.all_combinations = list(itertools.combinations(range(self.n), self.k))
15 self.subset_indices_tensor = torch.tensor([list(s) for s in self.all_combinations], device=self.device)
16 self.subset_excluded_tensor = torch.tensor(
17 [
18 [i for i in range(self.n) if i not in subset_indices]
19 for subset_indices in self.all_combinations
20 ], device=self.device
21 )
22 self.total_violations = len(self.all_combinations) * self.k * (self.n-self.k)
23
24 def calculate_loss(self):
25 # Convert batch_subset_indices to a tensor for efficient indexing
26 # Calculate sums of vectors for each subset in the batch
27 # Shape: (batch_size, d)
28 subset_sums = self.vector_embeddings[self.subset_indices_tensor].mean(dim=1)
29 # Calculate dot products of subset sums with all vectors
30 # Shape: (batch_size, n)
31 dot_products_all = torch.matmul(subset_sums, self.vector_embeddings.T)
32 # Get dot products with vectors in the subset by indexing dot_products_all
33 # Shape: (batch_size, k)
34 dot_products_xi = torch.gather(dot_products_all, 1, self.subset_indices_tensor)
35 # For each subset in the batch, calculate dot products with vectors NOT in the subset
36 # We can create a mask to zero out elements within the subset
37 # This is faster than iterating through the batch
38 # Shape: (batch_size, n-k)
39 dot_products_xj = torch.gather(dot_products_all, 1, self.subset_excluded_tensor)
40 # Calculate differences and apply ReLU
41 # We need to unsqueeze dot_products_xi to match dimensions for broadcasting
42 # Shape: (batch_size, n-k) - (batch_size, k)
43 differences = dot_products_xi.unsqueeze(1) - dot_products_xj.unsqueeze(2)
44 total_loss = torch.relu(-differences).sum((-1, -2)).mean()
45 num_violations = (differences < 0).sum().item()
46 return total_loss, num_violations
47
48
49 def train(self, d, num_epochs, learning_rate=0.1, patience=20):
50 self.vector_embeddings = torch.randn(
51 self.n, d, device=self.device, requires_grad=True
52 )
53
54
55 optimizer = optim.Adam(
56 [self.vector_embeddings],
57 lr=learning_rate,
58 )
59 scheduler = optim.lr_scheduler.OneCycleLR(
60 optimizer=optimizer,
61 max_lr=learning_rate,
62 total_steps=num_epochs,
63 pct_start=0.0,
64 )
65
66 min_violations = self.total_violations
67 epochs_no_improve = 0
68
[…]
70 for epoch in epoch_iterator:
71 optimizer.zero_grad()
72
73 loss, violations = self.calculate_loss()
74 loss.backward()
75 optimizer.step()
76 scheduler.step()
77
78
79 epoch_loss = loss.item()
80 epoch_iterator.set_postfix(
81 {
82 "n": self.n,
83 "k": self.k,
84 "loss": epoch_loss,
85 "#viorate": violations / self.total_violations,
86 "min#vio": min_violations,
87 "epochs_no_improve": epochs_no_improve
88 }
89 )
90
91 if violations < min_violations:
92 min_violations = violations
93 epochs_no_improve = 0
94 else:
95 epochs_no_improve += 1
96
97 if violations == 0:
98 print("Earlystopping:Noviolationsfound.")
99 break
100 if epochs_no_improve >= patience:
101 print(f"Earlystopping:Noimprovementinviolationsfor{patience}epochs.")
102 break
103
104 return min_violations
105
106
107class Experiment:
108 def __init__(self):
109 self.search_paths = {}
110 self.minimal_dimensions = {}
111
112 def find_minimal_dimension(self, k, n_values, left0=0, num_epochs=100, learning_rate=0.1, patience=10):
113
114 last_minimal = left0
115 for n in n_values:
116 print("#" * 10 + "newtask" + "#" * 10)
117 print(f"Findingminimaldimensionforn={n},k={k}")
118 print("#" * 30)
119 trainer = Trainer(n, k)
120
121 self.search_paths[n] = []
122 left, right = last_minimal+1, last_minimal + 40 # the range is set by prior to accelerate the convergence.
123 minimal_d = n + 1
124
125 while left <= right:
126 mid = (left + right) // 2
127 if mid == 0:
128 mid = 1
129
[…]
131
132 violations = trainer.train(
133 d=mid,
134 num_epochs=num_epochs,
135 learning_rate=learning_rate / np.log2(n),
136 patience=patience
137 )
138
[…]
140
141 self.search_paths[n].append({’dimension’: mid, ’violations’: violations})
142
143 if violations == 0:
144 minimal_d = mid
145 right = mid - 1
146 else:
147 left = mid + 1
148
149
150 self.minimal_dimensions[n] = minimal_d if minimal_d <= n else -1
151 last_minimal = minimal_d
152
153 with open("minimal_dem_log.txt", "at") as f:
[…]
155 print("#" * 10 + "TaskFinished" + "#" * 10)
[…]
157 print("#" * 30)
158
159 return self.minimal_dimensions
160
161 def plot_minimal_dimension_vs_n(self, k):
162 if not self.minimal_dimensions:
163 print("Noexperimentresultstoplot.Runfind_minimal_dimensionfirst.")
164 return
165
166 n_plot = list(self.minimal_dimensions.keys())
167 d_plot = list(self.minimal_dimensions.values())
168
169 plt.figure(figsize=(8, 6))
170 plt.plot(n_plot, d_plot, marker=’o’, linestyle=’-’)
171 plt.xlabel("NumberofVectors(n)")
172 plt.ylabel("MinimalDimension(d)")
173 plt.xscale(’log’)
174 plt.title(f"MinimalDimensionvs.NumberofVectorsfork={k}")
175 plt.grid(True)
176 plt.show()
177
178 plt.savefig("med_vs_n.png")
179
180 def plot_violations_vs_d(self, k):
181 if not self.search_paths:
182 print("Noexperimentresultstoplot.Runfind_minimal_dimensionfirst.")
183 return
184
185 for n, path in self.search_paths.items():
186 sorted_path = sorted(path, key=lambda x: x[’dimension’])
187
188 dimensions = [item[’dimension’] for item in sorted_path]
189 violations = [item[’violations’] for item in sorted_path]
190
191 plt.figure(figsize=(8, 6))
192 plt.plot(dimensions, violations, marker=’o’, linestyle=’-’)
193 plt.xlabel("Dimension(d)")
194 plt.ylabel("NumberofViolations")
195 plt.title(f"Violationsvs.Dimensionforn={n},k={k}")
196 plt.grid(True)
197 plt.show()
198
199
200# 1. Define a list of n_values to experiment with
201n_values = [5 * 2 ** (i) for i in [0, 1, 2, 3, 4, 5]]
202
203# 2. Set a value for k
204k = 2
205
206# 3. Instantiate the Experiment class
207experiment = Experiment()
208
209# 4. Call the find_minimal_dimension method with a reasonable batch_size and limited epochs
210# Using a batch_size of 32 and num_epochs of 100 for initial testing.
211print("Startingexperimentwithmini-batchingandearlystopping...")
212minimal_dims = experiment.find_minimal_dimension(
213 k, n_values, learning_rate=1, num_epochs=1000, patience=1000)
214
215# 5. Print the minimal_dims dictionary
[…]
217
218# 6. Call the plot_minimal_dimension_vs_n method
219experiment.plot_minimal_dimension_vs_n(k)
220
221# 7. Call the plot_violations_vs_d method
222experiment.plot_violations_vs_d(k)